OpenStudy (anonymous):

The cooking compartment of a microwave oven measures 15 inches long, 14 inches wide, and 9 inches high. What is the volume of the cooking compartment of this microwave oven? (will give out fan) :)

OpenStudy (anonymous):

V= L*W*H V=15*14*9 V=1890